OpenStudy (anonymous):

If the scale used for a model is 1:48, then 2 1/2 centimeters on the model represents 120 meters. True or False??? Please help me! I've spent at least an hour already!!!

jimthompson5910 (jim_thompson5910):

2 and 1/2 cm = 2 + 1/2 cm 2 and 1/2 cm = 2 + 0.5 cm 2 and 1/2 cm = 2.5 cm

jimthompson5910 (jim_thompson5910):

the ratio of scale:real is 1:48 ie, if you have 1 cm on the scale model, then you multiply it by 48 to get 48 cm so 1 cm corresponds to 48 cm It is certainly possible to replace the "cm" with "m" if you want the final real world units to be bigger So... 1 cm : 48 m 2.5*1 cm : 2.5*48 m 2.5 cm : 120 m So 2.5 cm on a scale model corresponds to 120 m on the actual object

OpenStudy (anonymous):

Thank you~! This made more sense ^_^

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (phi):

unless they explicitly say 1 cm : 48 m I would assume 1:48 means 1 cm represents 48 cm that means 2.5 cm represents 2.5*48 or 120 cm or 1.2 meters